OpenStudy (anonymous):

solve... cos^2 x+2cos x=0

OpenStudy (anonymous):

cosx{cos x + 2) = 0 cos x = 0 or cos x = -2 there is angle with cos = -2 x = 90 degrees or 20 degrees

OpenStudy (anonymous):

sorry - no angle with cos = -2 x = 90 or 270 degrees

OpenStudy (anonymous):

\[\cos^2 x + 2\cos x=0\]

OpenStudy (anonymous):

does this clear things up

OpenStudy (anonymous):

He already worked it out above for you. The question to you: does that clear things up?

OpenStudy (anonymous):

not really sorry

OpenStudy (anonymous):

You have to say what you don't understand; because he did it step for step.

OpenStudy (anonymous):

i dont understand what he wrote can you use the equation button

OpenStudy (anonymous):

i dont understand what he wrote can you use the equation button

OpenStudy (anonymous):

The equation button helps if there is an exponent or a base number. It is just writing an equation.

OpenStudy (anonymous):

how about we start from the begining

OpenStudy (anonymous):

All he did was 'pull' the cos x out. For example, suppose it was 4+2 That is the same thing as 2(2+1) That is what he did to your cos x.

OpenStudy (anonymous):

oh ok now i get it